Express the function 𝑓(π‘₯) = π‘₯ /(π‘₯βˆ’3)2 as partial fractions and hence find ∫ 𝑓(π‘₯)𝑑x


1
Expert's answer
2021-06-07T13:32:44-0400

Solution

The function 𝑓(π‘₯) = π‘₯ /(π‘₯βˆ’3)2Β may be represented as sum of partial fractions f(x) = A/(x-3)+B/(x-3)2Β where A, B are arbitrary constants. From equality π‘₯ /(π‘₯βˆ’3)2 = A/(x-3)+B/(x-3)2 we’ll get Β 

A(x-3)+B = xΒ => A=1, B=3A=3 => 𝑓(π‘₯) = 1/(π‘₯βˆ’3)+3/(π‘₯βˆ’3)2 Β Β Β 

Now ∫ 𝑓(π‘₯)𝑑x = ∫ [1/(π‘₯βˆ’3)+3/(π‘₯βˆ’3)2]𝑑x = ∫ [1/(π‘₯βˆ’3)]𝑑x + 3∫ [1/(π‘₯βˆ’3)2]𝑑x = ln|x-3|-3/(π‘₯βˆ’3)+CΒ Β 

Answer

𝑓(π‘₯) = 1/(π‘₯βˆ’3)+3/(π‘₯βˆ’3)2

∫ 𝑓(π‘₯)𝑑x = ln|x-3|-3/(π‘₯βˆ’3)+C